I'm trying to help my 7th grader with an algebra problem. Boy...I feel stupid! Here it is.
A bike shop has a bunch of bikes for sale. The problem asks how many of each bike are available. Here is the information given:
The shop sells tricycles, bicycles, tandem bikes, and unicycles.
There are 57 helmets available (one for each rider).
There are 115 wheels total.
There are 32 regular bikes.
There are the same same number of tandem bikes as ther are unicycles.
That is all the information there is. Thanks for your help!
Rich

Answer by mananth(16946)

t= tricycle
b-bicycle
d= tandem
u- unicycle
There are 57 helmets available (one for each rider).
t+b+2d+u=57
There are 115 wheels total.
3t+2b+2d+u=115

There are 32 regular bikes.
b=32

There are the same same number of tandem bikes as there are unicycles.
d=u
t+b+2d+u=57
3t+2b+2d+u=115
b=32
d=u
Substitute in the first equation
t+32+2d+d=57
t+3d=25>...............................(1)
substitute in second equation
3t+64+2d+d=115
3t+3d=115-64
3t+3d=51
t+d=17..................................(2)
solve 1 & 2 we get
d= 4 Tandem bikes
t+d=17
t=13.... Tricycles
u = d
so u =4